OpenStudy (anonymous):

Surds question! Simplify (7 + 3 x squareroot2)^2 - (5 - 2 x squareroot5)^2

OpenStudy (nubeer):

hint (a+b)^2 = a^2 +2ab+b^2 (a-b)^2 = a^2 -2ab +b^2

OpenStudy (anonymous):

 Is this correct?